### v1

@@ Abstract
Multi-agent systems built on large language models (LLMs) promise enhanced problem-solving capabilities through collaborative division of labor. However, they frequently underperform single-agent baselines due to edge-level error cascades: minor inaccuracies at one message handoff propagate across the entire chain. We propose AgentAsk, a lightweight and plug-and-play clarification module that treats every inter-agent message as a potential failure point and inserts minimally necessary questions to arrest error propagation. AgentAsk follows a three-stage pipeline: (i) distilling edge-level judgments from curated failure traces into a compact policy, (ii) supervising the policy to determine when/what/whom/how to ask, and (iii) optimizing online with E-GRPO, a reinforcement learning objective that balances accuracy, latency, and cost. The module is architecture-agnostic and easy to integrate into existing orchestration. Across math, reasoning, and coding benchmarks, AgentAsk consistently improves accuracy and robustness over public multi-agent implementations while keeping overhead minimal, with latency and extra cost all less than 5%, approaching the performance of a strong evaluator. Beyond empirical improvements, we contribute a principled taxonomy of edge-level errors and a practical recipe for link-local intervention, offering a scalable pathway toward more reliable LLM-based multi-agent systems. The code is available at https://github.com/BoHan-LIN04/AgentAsk.
@@ 1 Introduction
Large language model (LLM)–based agent systems have attracted increasing attention for their ability to combine multi-step reasoning with dynamic tool use (Wei et al., 2022a; Schick et al., 2023). By coordinating multiple LLM-driven agents, these systems aim to leverage collective intelligence to solve complex, real-world tasks that necessitate decomposition, external knowledge access, and iterative refinement(Park et al., 2023). This paradigm is referred to as a multi-agent system (MAS). Applications now span diverse domains such as software engineering (Hong et al., 2024; Qian et al., 2024) and AI for Science (Ghafarollahi and Buehler, 2025; M. Bran et al., 2024), and there is growing interest in developing general-purpose multi-agent frameworks for cross-domain problem solving (Wu et al., 2024a; Liu et al., 2024; Yu et al., 2025a). These advances underscore both the promise of multi-agent collaboration and the difficulty of making MAS reliable in practice.
Figure 1: Multi-Agent System Reasoning: Success vs. Cascade Failure. The left shows the normal operation of MAS, while the right shows that the upstream agents transmit errors during the interaction that trigger a cascade effect and cause the entire system to fail.
Despite the promise of collaborative decomposition, multi-agent systems often fail to reliably surpass strong single-agent baselines across diverse tasks. Previous research and benchmarks report unstable gains and brittle behavior under realistic orchestration (Liu et al., 2024; Cemri et al., 2025). Execution logs reveal a recurring pattern in which small inconsistencies or missing details introduced at one step travel along the chain and accumulate into system-level failures. Building on this pattern, chainwise amplification is intensified by specification gaps and inter-agent misalignment, and in some settings, collaboration reduces accuracy (Zhang et al., 2025d; Wynn et al., 2025; Wang et al., 2025). Prior studies quantify the effect, with success on difficult programming tasks falling to about 25% (Cemri et al., 2025). Similar vulnerabilities appear in single-agent traces, including misinterpretations, logical gaps, and limited reflection, which indicates that subtle errors arise early and persist through execution. This sets up the problem we study: limiting error growth at the handoff between agents so that small inconsistencies do not accumulate into system-level failures.
@@
We argue that ensuring reliability necessitates a paradigm shift toward edge-level intervention: each inter-agent message should be treated as a potential failure point where minimal clarification can halt cascading errors, illustrated in Figure 1. To this end, we propose AgentAsk, a plug-and-play clarification module for LLM-based multi-agent systems. AgentAsk introduces a taxonomy of four core edge-level error types and implements a three-stage pipeline: (i) distilling knowledge from failure traces with a powerful evaluator, (ii) transferring this distilled capability to a lightweight clarifier trained to generate targeted questions, and (iii) refining the clarifier through reinforcement learning for adaptive, budget-aware behavior. Crucially, AgentAsk is designed to be lightweight, architecture-agnostic, and easily integrated into existing frameworks, enabling systems to intercept misunderstandings before they spread.
@@
  • Edge-level perspective and taxonomy: We formalize a four-type taxonomy of inter-agent errors and motivate clarification as a first-class mechanism for multi-agent collaboration.
  • Practical clarification module: AgentAsk is a lightweight and plug-and-play module trained with E-GRPO reinforcement learning method, which agents can seamlessly integrate to identify and resolve interaction errors.
  • Comprehensive empirical evaluation: Extensive experiments across diverse benchmarks demonstrate that AgentAsk consistently improves both accuracy and robustness over public multi-agent baselines with excessive latency and extra cost of only less than 5%.
@@ 2 Related Work
MAS coordinates multiple role-specialized agents, often equipped with tools, to solve complex tasks. Early systems demonstrated that organizing multiple LLMs into role-specialized, conversational teams with tool use can outperform single models under controlled settings, through frameworks that scaffold roles, protocols, and conversation programming such as AutoGen (Wu et al., 2024a), AgentVerse (Chen et al., 2024b), CAMEL (Li et al., 2023), MetaGPT (Hong et al., 2024), and ChatDev Qian et al. (2024). Beyond hand-crafted blueprints, recent work treats orchestration as a learnable object: GPTSwarm models agents and communication links as an optimizable graph (Zhuge et al., 2024), AFlow searches over code-represented workflows via Monte Carlo tree search (Zhang et al., 2025b), MaAS samples query-dependent sub-architectures from an agentic supernet (Zhang et al., 2025a), and MasRouter learns a cascaded controller for collaboration mode, role allocation, and LLM selection (Yue et al., 2025). These directions build on and complement decision-time tool use, such as ReAct (Yao et al., 2023) and Toolformer (Schick et al., 2023). Our work is orthogonal to global orchestration search: we hold an existing pipeline fixed and inject a lightweight edge-level clarifier that adds minimal, well-placed questions at message handoffs to curb error propagation with low latency and cost.
@@
Large-scale audits show that multi-agent gains over strong single-agent baselines are not guaranteed and that many failures arise at message handoffs where missing details, ambiguous references, or distorted intermediate results are left unclarified (Cemri et al., 2025; Zhu et al., 2025; Zhang et al., 2025d). Methodologically related attempts to improve reliability include multi-agent debate and discussion (Chan et al., 2024; Liang et al., 2024), self-feedback loops such as Reflexion Shinn et al. (2023) and Self-Refine (Madaan et al., 2023), and verification or self-correction procedures (Wu et al., 2024b; Lee et al., 2025), yet recent analyses document conditions where debate or naive self-correction can fail or even degrade accuracy (Wynn et al., 2025). A complementary line advocates asking clarifying questions under uncertainty in single-agent human–computer interaction(Wang et al., 2024; Zhang et al., 2025c; Mukherjee et al., 2025; Lee et al., 2023). We follow this clarification-centric view but operationalize it for multi-agent settings with an edge-aware policy that decides when, what, whom, and how to ask exactly at the handoff where uncertainty or distortion originates.
@@ 3 Error Taxonomy
Many studies catalog errors in multi-agent systems and report large-scale statistics, yet most emphasize breadth over depth. Prior classifications cover many error types but give limited attention to isolating the most critical ones or to prescribing targeted remedies. In particular, although prior classifications cover a wide range of error types, little attention has been paid to identifying the most critical errors and developing targeted remedies. In this work, we empirically reveal chain-style error propagation as a fundamental root cause of failures in multi-agent systems, where a single error of an agent can cascade into system-wide collapse. We audit 824 execution logs and annotate each agent-to-agent message as the unit of analysis, building on the existing research (Wang et al., 2024). Cross annotation and adjudication yield a concise taxonomy with high agreement that concentrates intervention exactly where failures begin.
Analogous to human collaboration, our empirical results (shown in Figure 2) indicate that most errors in MAS stem from violations of a fundamental principle: agents should complete tasks accurately within their capabilities and deliver clear and complete information handovers to downstream agents without omission. Building on this observation, four types of errors have been condensed.
@@ 3.4 Capability Gap
Figure 3: Overview of AgentAsk. The module operates at the edge level, treating each inter-agent message as a potential failure point. The figure illustrates both the architecture and training process: (i) knowledge distillation from failure traces using a large evaluator to construct an edge-level corpus, (ii) supervised fine-tuning of a lightweight clarifier that decides when/what/whom/how to ask, and (iii) reinforcement learning with E-GRPO for adaptive clarification under latency and cost constraints. This pipeline equips AgentAsk with edge-aware monitoring and minimal, targeted interventions while remaining architecture-agnostic and easy to integrate into diverse MASs.
3.5 Inspiration
Our analysis shows that all four error types reduce to failures at the edge, the information handoff between agents in a multi-agent system. This shifts the design focus to the handoff itself rather than the global workflow. In our annotated corpus NN=824, these errors account for most failures, with Data Gap 29.1%, Referential Drift 27.3%, Signal Corruption 36.8%, and Capability Gap 6.8%. In effective human collaboration (Woolley et al., 2010; Cobbe et al., 2021), handoffs are governed by three habits: ❶ check the handoff for missing content, ambiguous reference, corrupted intermediate results, or role mismatch; ❷ if a problem exists, fix it at the source so it does not propagate downstream; ❸ intervene only as needed, keeping the exchange brief and within latency and cost limits.
These habits motivate an edge-centered design that monitors each transfer and issues minimal clarifications exactly at the point of exchange. The details can be seen in Appendix 9.
@@ 4 Design of AgentAsk
Grounded in the taxonomy, failures cluster at the edge where one agent hands information to the next. We therefore design around three moves and introduce AgentAsk, a framework for real-time monitoring and correction in multi-agent systems. First, to find errors, we place an external clarifier at each handoff that inspects the outgoing message and its local context and flags risk early. Second, to correct errors, we use “Ask”: the AgentAsk sends a brief, targeted question to the right party so the handoff is repaired before the chain advances. Third, we transferred the asking capabilities of the powerful model to AgentAsk, but to keep correction efficient, fine-tuning alone is not enough, so we add reinforcement learning that adapts the ask decision, the addressee, and the phrasing to context and budget. Concretely, we use a two-stage training paradigm, Knowledge Transfer via Supervised Fine-tuning followed by Autonomous Learning via Reinforcement Optimization, which gives AgentAsk an inherited ability to detect faulty outputs and a learned ability to apply minimal, cost-aware clarifications. The empirical experiments are provided in Section 5. The design of the proposed method is illustrated in Figure 3. Complete training pseudocode appears in Appendix A as Algorithm 1.
4.1 Modeling of Multi-Agent System
We view a multi-agent system as a directed interaction graph $G=(V,E)$ that is unrolled per input instance into an edge sequence $\{e_{t}\}_{t=1}^{T}$, where each edge $e_{t}=(u_{t}\!\rightarrow\!v_{t})$ carries a message produced by agent $u_{t}$ and consumed by agent $v_{t}$. AgentAsk is an edge-local controller that sits between $u_{t}$ and $v_{t}$ and decides whether to intervene with a minimal clarification before the message is delivered; the decision must respect latency and budget while improving task success.
We represent the edge context by a compact state that contains only what is available at the handoff:
$x_{t}=(x^{\mathrm{in}}_{t},\,u_{t},\,v_{t},\,m_{t},\,h_{t}).$ (1)
Here $x^{\mathrm{in}}_{t}$ denotes the user query or current subgoal, $m_{t}$ is the candidate message from $u_{t}$, and $h_{t}$ encodes short-horizon interaction traces relevant to this link . The controller produces a structured action $a_{t}=(z_{t},\tilde{v}_{t},q_{t})$, where $z_{t}\in\{0,1\}$ is an ask gate, $\tilde{v}_{t}\in\{u_{t},v_{t}\}$ is the addressee when asking, and $q_{t}$ is a short clarification string following a fixed schema.
The decision policy is factored to separate the binary gate, the addressee, and the question generator while falling back to no-op when $z_{t}=0$:
$\pi_{\theta}(a_{t}\mid x_{t})=\pi_{\theta}(z_{t}\mid x_{t})\,\Big[\pi_{\theta}(\tilde{v}_{t}\mid x_{t})\,\pi_{\theta}(q_{t}\mid x_{t})\Big]^{z_{t}}.$ (2)
This factorization treats the clarification as conditional text generation under a hard budget on length and a lightweight schema constraint; when $z_{t}=0$, the message passes through unchanged, and the edge terminates immediately.
After executing $a_{t}$, the environment returns a reply $r_{t}$ when $z_{t}=1$ and updates the local trace and counters, leading to the next edge state $x_{t+1}$ and a deterministic delivery to $v_{t}$. The full interaction for one input forms a trajectory $\tau=\{(x_{t},a_{t})\}_{t=1}^{T}$ with a terminal outcome that scores correctness, while the controller accrues latency and tokenized cost along the way; we summarize this as a constrained objective over trajectories
$\max_{\theta}\ \mathbb{E}_{\tau\sim\pi_{\theta}}[U(\tau)]\ \ \text{s.t.}\ \ \mathbb{E}_{\tau\sim\pi_{\theta}}[C(\tau)]\leq B.$ (3)
Here $U(\tau)$ is a task utility that rewards accurate solutions and well-formed clarifications, $C(\tau)$ aggregates latency and token usage induced by interventions, and $B$ is a budget. Within this modeling view, the next two stages instantiate the policy $\pi_{\theta}$: a supervised stage that transfers edge-level judgment and a reinforcement stage that optimizes the ask gate and question under the constraint in Eq. (3).
4.2 Knowledge Transfer via Supervised Fine-tuning
We initialize the edge controller by supervised learning to imitate high-quality clarifications on individual handoffs.
Corpus construction.
The SFT corpus is built from logged multi-agent traces by applying a teacher judge that assigns, for each edge state $x_{i}$ from Eq. (1), a minimal intervention triple $y_{i}=(t_{i},v_{i},q_{i})$ with $t_{i}\in\mathcal{T}$, $v_{i}\in\{u_{i},v_{i}\}$, and a short question $q_{i}$. Concretely, we construct
$\widetilde{\mathcal{D}}$ $=\{(x_{i},y_{i})\}_{i=1}^{N},\ \ y_{i}=(t_{i},v_{i},q_{i}),$ (4)
The label $t_{i}$ encodes the ask decision and type, $v_{i}$ is the addressee when asking, and $q_{i}=(q_{i,1},\ldots,q_{i,T_{i}})$ is a schema-constrained question with a token cap; the indicator $m_{i}=\mathbbm{1}[t_{i}\neq\texttt{NONE}]$ gates whether an ask is present. We train the factored policy in Eq. (2) with three heads: a classifier $p_{\theta}^{\text{type}}(t\mid x)$, an addressee head $p_{\theta}^{\text{addr}}(v\mid x,t)$, and an auto-regressive decoder $p_{\theta}^{\text{txt}}(q\mid x,t)$.
Supervised transfer.
The SFT objective separates the gate/type supervision from the ask-conditioned components and then combines them. The type loss is
The final SFT objective is a weighted sum that matches the factorization and respects the gate:
$\mathcal{L}_{\text{SFT}}=\mathcal{L}_{\text{type}}+\lambda_{\text{ask}}\mathcal{L}_{\text{ask}}.$ (6)
This trains the controller to first decide whether and what to ask, and, only when needed, to select the recipient and generate a concise, well-formed question aligned with edge-local budgets. In particular, CG questions can be solved to some extent by the target address when asking the question.
4.3 Autonomous Learning via Reinforcement Optimization

### v2

@@ Abstract
Multi-agent systems (MAS) built on large language models promise improved problem-solving through collaboration, yet they often fail to consistently outperform strong single-agent baselines due to error propagation at inter-agent message handoffs. In this work, we conduct a systematic empirical analysis of such failures and introduce an edge-level error taxonomy that identifies four dominant error types: Data Gap, Signal Corruption, Referential Drift, and Capability Gap, as primary sources of failure in multi-agent interactions. Building on this taxonomy, we propose AgentAsk, a lightweight clarification module designed to intervene at the edge level in MAS to prevent cascading errors. The module operates by strategically applying minimal clarifications at critical points within the system, improving the accuracy and efficiency of the overall task. AgentAsk is trained to balance the trade-offs between clarification cost, latency, and accuracy, while it is also architecture-agnostic and can be easily integrated into existing systems. Evaluated across five benchmarks, AgentAsk consistently improves accuracy by up to 4.69%, while keeping latency and extra costs below 10% compared to baseline MAS, showcasing its high efficiency and minimal overhead.
@@ 1 Introduction
Large language model (LLM)–based agent systems have garnered attention due to their ability to perform multi-step reasoning and dynamic tools. By coordinating multiple LLM-driven agents, these systems leverage collective intelligence to tackle complex tasks that require decomposition, external knowledge access, and iterative refinement (Park et al., 2023). This approach, known as multi-agent systems (MAS), has the potential to solve a wide range of real-world problems (Liu et al., 2024; Yu et al., 2025a; Wu et al., 2024a).
Figure 1: Multi-Agent System Reasoning: Success vs. Cascade Failure. The left illustrates successful multi-agent reasoning with consistent information exchange, while the right shows how errors introduced by upstream agents propagate across interactions, triggering a cascade that leads to system-level failure.
Despite the promise of collaborative decomposition, multi-agent systems often fail to reliably surpass strong single-agent baselines across diverse tasks. Previous research and benchmarks report unstable performance gains and brittle behavior under realistic orchestration (Liu et al., 2024; Cemri et al., 2025). Execution logs reveal a recurring pattern in which small inconsistencies or missing details introduced at one step travel along the chain and accumulate into system-level failures. Figure 1 illustrates this process. Building on this, chainwise amplification is intensified by specification gaps and inter-agent misalignment, and in some settings, collaboration reduces accuracy (Zhang et al., 2025d; Wynn et al., 2025; Wang et al., 2025). Prior studies quantify the effect, with success on difficult programming tasks falling to about 25% (Cemri et al., 2025). Similar vulnerabilities appear in single-agent traces, including misinterpretations, logical gaps, and limited reflection, which indicates that subtle errors arise early and persist through execution. These observations motivate the problem we study: how to limit error growth at inter-agent handoffs, such that small local inconsistencies do not accumulate into system-level failures.
@@
To this end, we argue that ensuring reliability necessitates a paradigm shift toward edge-level intervention: each inter-agent message should be treated as a potential failure point where minimal clarification can halt cascading errors. Based on this perspective, we introduce a taxonomy of four core edge-level error types that are responsible for most failures in MAS interactions. Guided by these insights, we propose AgentAsk, a plug-and-play clarification module for LLM-based multi-agent systems. AgentAsk operates by identifying potential failure points at each message handoff and applying minimal clarification to prevent cascading errors. Crucially, AgentAsk is designed to be lightweight, architecture-agnostic, and easily integrated into existing frameworks, enabling systems to intercept misunderstandings before they spread.
@@
  • Edge-level perspective and taxonomy: We introduce an edge-level perspective on multi-agent failures and formalize a taxonomy of four inter-agent error types, motivating clarification as an effective mechanism for improving multi-agent collaboration.
  • Practical clarification module: We propose AgentAsk, a lightweight and plug-and-play clarification module that uses the edge-level taxonomy to identify and resolve interaction errors at message handoffs, improving system reliability with minimal intervention.
  • Comprehensive empirical evaluation: Extensive experiments across multiple benchmarks show that AgentAsk consistently enhances accuracy by up to 4.69%, while keeping latency and extra cost below 10% compared to existing multi-agent baselines.
@@ 2 Related Work
MAS coordinates multiple role-specialized agents, often equipped with tools, to solve complex tasks. Early systems demonstrated that organizing multiple LLMs into role-specialized, conversational teams with tool use can outperform single models, through frameworks that scaffold roles, protocols, and conversation programming such as AutoGen (Wu et al., 2024a), AgentVerse (Chen et al., 2024b), CAMEL (Li et al., 2023), MetaGPT (Hong et al., 2024), and ChatDev Qian et al. (2024). Recent work treats orchestration as a learnable object: GPTSwarm models agents and communication links as an optimizable graph (Zhuge et al., 2024), AFlow searches over code-represented workflows via Monte Carlo tree search (Zhang et al., 2025b), MaAS samples query-dependent sub-architectures from an agentic supernet (Zhang et al., 2025a), and MasRouter learns a cascaded controller for collaboration mode, role allocation, and LLM selection (Yue et al., 2025). These directions build on and complement decision-time tool use, such as ReAct (Yao et al., 2023) and Toolformer (Schick et al., 2023). Our work is orthogonal to global orchestration search: we hold an existing pipeline fixed and inject a lightweight edge-level clarifier that adds questions at message handoffs to curb error propagation.
@@
Large-scale audits show that multi-agent gains over strong single-agent baselines are not guaranteed and that many failures arise at message handoffs(Cemri et al., 2025; Zhu et al., 2025; Zhang et al., 2025d). Methodologically related attempts to improve reliability include multi-agent debate and discussion (Chan et al., 2024; Liang et al., 2024), self-feedback loops such as Reflexion Shinn et al. (2023) and Self-Refine (Madaan et al., 2023), and verification or self-correction procedures (Wu et al., 2024b; Lee et al., 2025), yet recent analyses document conditions where debate or naive self-correction can fail or even degrade accuracy (Wynn et al., 2025). A complementary line advocates asking clarifying questions under uncertainty in single-agent human–computer interaction(Wang et al., 2024; Zhang et al., 2025c; Mukherjee et al., 2025; Lee et al., 2023). We follow this clarification-centric view but operationalize it for multi-agent settings with an edge-aware policy that decides when, what, whom, and how to ask exactly at the handoff.
@@ 3 Error Taxonomy
Conceptual Inspiration. In multi-agent systems, ensuring reliable communication between agents is essential to prevent cascading errors that can propagate throughout the system. Our error taxonomy is inspired by human collaboration theory (Gittell, 2006; Woolley et al., 2010), including the concept of Relational Coordination, which emphasizes the importance of frequent, timely, and accurate communication, shared goals, and mutual respect among team members. According to this theory, coordination can falter when communication is unclear, roles are misaligned, or team members fail to anticipate each other’s needs, leading to errors in the system.
Empirical Evidence. In this work, we empirically demonstrate that chain-style error propagation is a fundamental root cause of failures in MAS, where a single error can cascade into a system-wide collapse. We reveal these errors through an audit of 824 execution logs, annotating each agent-to-agent message as the unit of analysis. To ensure the reliability of our classification, the annotation process was conducted by multiple professional annotators with expertise in MAS. These annotators engaged in several rounds of cross-annotation and adjudication to achieve consistent results. The high inter-rater reliability was confirmed with a Fleiss’ Kappa value of 0.84, indicating substantial agreement among annotators. Our empirical results, shown in Figure 2, indicate that most errors in MAS stem from violations of a fundamental principle: agents should complete tasks accurately within their capabilities and deliver clear and complete information handovers to downstream agents without omission. Building on this observation, four types of errors have been condensed.
@@ 3.4 Capability Gap
Figure 3: Overview of AgentAsk. The module operates at the edge level, treating each inter-agent message as a potential failure point. The figure illustrates both the architecture and training process: (i) Constructing an edge-level corpus from failure logs by a large evaluator, (ii) Learning to ask from demonstrations to decides when/what/whom/how to ask, and (iii) reinforcement learning with E-GRPO for clarification optimization under latency and cost constraints. This pipeline equips AgentAsk with edge-aware monitoring and minimal, targeted interventions while remaining architecture-agnostic and easy to integrate into diverse MASs.
3.5 Inspiration from Edge-level Taxonomy
Our analysis shows that all four error types manifest as failures at the edge, the information handoff between agents in a multi-agent system. This shifts the design focus to the handoff itself rather than the global workflow. In our annotated corpus (NN=824), these errors account for most failures, with Data Gap 29.1%, Referential Drift 27.3%, Signal Corruption 36.8%, and Capability Gap 6.8%. According to the human collaboration theory we mentioned above, effective handoffs are governed by three habits: ❶ check the handoff for missing content, ambiguous reference, corrupted intermediate results, or role mismatch; ❷ if a problem exists, fix it at the source so it does not propagate downstream; ❸ intervene only as needed, keeping the exchange brief and within latency and cost limits.
These habits motivate an edge-centered design that monitors each transfer and issues minimal clarifications exactly at the point of exchange. The illustrative examples can be seen in Appendix D.
@@ 4 Design of AgentAsk
Overview. Grounded in the taxonomy, failures cluster at the edge where one agent hands information to the next. We therefore design around three moves and introduce AgentAsk, a framework for real-time monitoring and correction in multi-agent systems. First, to identify errors, we place an external clarifier at each handoff that inspects the outgoing message and its local context, flagging risks early. Second, to correct errors, we use “Ask”: the AgentAsk sends a brief, targeted question to the right party so the handoff is repaired before the chain advances. Third, we transferred the asking capabilities of the powerful model to AgentAsk, but to keep correction efficient, fine-tuning alone is not enough, so we add reinforcement learning that adapts the ask decision, the addressee, and the phrasing to context and budget.
The training process of the framework is detailed in the following chapters, while its overall design is outlined in Figure 3. Complete training pseudocode appears in Appendix B as Algorithm 1, and the notation used throughout is summarized in Table 14 of Appendix E.
4.1 Clarification as Edge Intervention
We model multi-agent collaboration as a sequence of message handoffs over a directed interaction graph. At each step, the controller AgentAsk receives a local view of the exchange and decides whether to insert a clarification to improve task outcomes. Each decision is based on an edge-local state $x_{t}$, where $t$ is the timestep of the handoff, and results in an action $a_{t}=(z_{t},\tilde{v}_{t},q_{t})$, where $z_{t}$ is a binary gate indicating whether to ask, $\tilde{v}_{t}$ selects which agent to query, and $q_{t}$ is the generated clarification question. The controller either allows the message to proceed or poses a brief clarification to one of the agents. This intervention is lightweight but can substantially improve task-level accuracy.
The policy $\pi_{\theta}$ governs the action sequence across an input, yielding a trajectory $\tau=\{(x_{t},a_{t})\}_{t=1}^{T}$. The controller is optimized to maximize expected task utility while operating under a cost constraint:
$\max_{\theta}\;\mathbb{E}_{\tau\sim\pi_{\theta}}\left[U(\tau)\right]\quad\text{s.t.}\quad\mathbb{E}\left[C(\tau)\right]\leq B.$ (1)
Here, $U$ is the utility function measuring task success, and $C$ represents the cost incurred from clarifications, while $B$ is the predefined budget limit. The internal structure of $x_{t}$ and $a_{t}$, as well as the factorized form of $\pi_{\theta}$, are introduced and discussed in Appendix A.1.
4.2 Learning to Ask from Demonstrations
We initialize the controller under a cold-start condition using fine-tuning on curated clarification traces. This allows AgentAsk to acquire structured ask behavior before reinforcement optimization. We construct a dataset of edge-local examples, each consisting of a context $x_{i}$ and an intervention $y_{i}$. Specifically, $x_{i}$ represents the state at edge $i$, and $y_{i}$ encodes the teacher’s decision (whether to ask, whom to ask, and what to ask). The controller learns to predict the same content.
Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ be the training dataset, where $N$ is the total number of examples. The training loss includes a decision component $\mathcal{L}_{\text{type}}$ for predicting whether a clarification is needed, and a conditional ask component $\mathcal{L}_{\text{ask}}$ for predicting the clarification details:
$\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{type}}+\lambda_{\text{ask}}\,\mathcal{L}_{\text{ask}}.$ (2)
Here, $\lambda_{\text{ask}}$ is a parameter that balances two components. This provides a cold-start policy capable of clarification. Details of the data construction, label format, and loss terms appear in Appendix A.2.
4.3 Clarification Optimization with Reinforcement Learning